OpenStudy (anonymous):

Set as a system of equation then solve: It takes a boat 1.5 hours to go 24 miles downstream and 3 hours to return upstream to its starting point. What is the boat's rate in still water? Please help me

OpenStudy (anonymous):

we need a variable for the rate in still water what do you pick?

OpenStudy (anonymous):

y ?

OpenStudy (anonymous):

or better still lets find the rate going against the current, and the rate going with the current that is easier, since you know both

OpenStudy (anonymous):

okay so will it be 1.5x+3y=24?

OpenStudy (danjs):

sorry, i was doing something else, didn't get back to the other one you posted

OpenStudy (anonymous):

distance is rate times time, and rate is distance divide by time, so the rate going downstream is \[\frac{24}{1.5}=16\] and the rate going upstream is \[\frac{24}{3}=8\]

OpenStudy (anonymous):

take the average of those to to get the rate in still water

OpenStudy (danjs):

d = r * t same distance down and up rate down * time down = rate up * time up when the water is moving also in same direction as boat, the relative rates are different

OpenStudy (danjs):

vs land observer speed is water + boat and water - boat with or against current

OpenStudy (anonymous):

so since we don't have the rate for upstream will it be 3x=1.5*24????

OpenStudy (danjs):

you have both distance and time for each trip, above shown rate is 16 downstream and 8 upstream

OpenStudy (danjs):

the velocity of the water and the velocity of the boat , are both relative, with respect to , a land spot, where the distance is 24

OpenStudy (anonymous):

Then do I have to plug in the rate of the boat going upstream and downstream to that?

OpenStudy (danjs):

oh , the system of equations thing you mentiond before is this

OpenStudy (danjs):

Given the distance and time down and back, you get the net rate in both directions those rates are a combo of both water and boat speeds,

OpenStudy (danjs):

the overall velocity down is 16 which is both the water and the boat velocities, back the overall rate is 8 and that is the water - boat velocities 16 = W + B 8 = W - B W = 12 B = 4